A disk pulley system is set up as shown. (Use moment of inertia of the center of a disk)r2 m2 TU

a. Symbolically find the value m1/m2 that would result in a net zero torque on the disks.

b. In the situation from part a, symbolically find the angular acceleration of the disks and tension in the ropes if m2 gains a new mass, M.
Answer #1

Solution in the uploaded images

Solution mass of pully = mp (a) 5 Torque = Forcex Idistance va nal m2 Net torque, Theta om, my Tnet = m_gx21 - miger mag m.geT2 = Ma Q = mg 2rxa [2Mom M 1 g ampa (2M-mi) ampa T2 = Mg- T2 = Mg [ g[24 Ti = mia + mig To Ti= mi g[2M-mi]) +mig (2M-mi) IT